Oil leaks from a tank. At hour there are 320 gallons of oil in the tank. Each hour after that, 4% of the oil leaks out. (a) What percent of the original 320 gallons has leaked out after 11 hours? % (b) If is the quantity of oil remaining after hours, find the value of .
Answer #1
A_1 = 320, r = 0.93

(a) 0.93^11 = 0.4501035457...

1 - 0.45 = 0.55

55% has leaked out after 11 hr.

(b) Q = Qo e^(kt)
0.55 = e^(k • 11)
0.55 = e^(11k)
-0.0543659348... = k

k ˜ -0.54 = -54%
Answer #2
oil leaked=0.04*(1+0.96+0.92...0.6)320=75.776 gallons oil left=320-75.776=244.224
Answer #3
solution
A_1 = 320, r =0.93
(a)
0.93^11 =0.4501035457...
1 - 0.45 = 0.55
55% has leaked out after 11 hr.



(b)
Q = Qo e^(kt)
0.55=e^(k • 11)
0.55=e^(11k)
-0.0543659348... = k

k ˜ -0.54 =-54%